# José Luis Blanco-Claraco

### v1

@@ 0.81 $d/hm$
MOLA-LO + LC
@@ 4.1 MulRan dataset
This multimodal dataset comprises Ouster OS1-64 LiDAR data, consumer-grade GNSS, and precise ground truth for 12 driving sequences (Kim et al. (2020)). All sequences include several loop closures, except the longest ones (Sejong{01,02,03}) which only include one, although the longest loop in all the benchmarked datasets ($>20$ km).
We have compared our LO with its default configuration against two other state-of-the-art LO methods: KISS-ICP (Vizzo et al. (2023)) using its default settings, and SiMpLE (Bhandari et al. (2024)) using their parameter set optimized for this dataset. MOLA wrappers have been provided for both methods to provide a fair comparison in terms of identical API to provide input raw data and to collect estimated trajectories. The MOLA dataset source (see §3.2) for MulRan takes care of removing LiDAR scans at the beginning or end of each sequence that have no associated ground truth. In this way, we can use the KITTI evaluation metrics (RTE and RRE) apart of APE (see metrics definitions in §4). Consumer-grade GNSS observations are not used by our LO system, but they are while applying the loop-closure post-processing stage (explained in §3.13).
@@
Figure 13: Estimated trajectories for the proposed LiDAR odometry system (”MOLA-LO”) applied to the KITTI odometry dataset, compared to ground truth. Trajectories denoted as “MOLA-LO + LC” include loop-closure. See discussion in §4.2.
Sample trajectories from our method are shown in Figure 12, compared to GT. It can be seen that the output from loop closure is nearly indistinguishable from GT. A quantitative benchmark is provided in Table 2, where all three metrics (RTE, RRE, APE) are provided, together with average run times per LiDAR scan (measured on an Intel i7-8700 at 3.20GHz). All LO methods are well below the sensor rate (10 Hz), hence all of them would be able to run onboard in real-time. An interesting observation is that there is not always a best performing method for a given sequence, since different methods may have the lowest values for one given metric only: LO methods would rank in a different order depending on the metric of interest. That said, for this dataset we can see that SiMpLE clearly outperforms the others, with our system coming out in second position. Hence, SiMpLE’s special optimization function is worth further investigation for its apparent superior robustness. Rankings apart, all three methods perform well in this dataset with the small performance differences not being probably relevant in practical robotics applications, where a full SLAM solution is normally required. Our SLAM version (reflected as “MOLA LO+LC” in the table) achieves, obviously, much better results than the pure LO methods, hence their values are not marked in bold since the method is qualitatively different.
@@ 4.2 KITTI-odometry dataset
The KITTI odometry dataset (Geiger et al. (2013)) was the first widely spread benchmark for LiDAR and visual odometry and SLAM methods, hence its popularity and importance. It comprises 11 training sequences (numbered 00-10) with public ground truth and other 11 evaluation sequences with undisclosed ground truth trajectories. Available sensors in this driving dataset include 3D LiDAR (a Velodyne HDL-64E) and two pairs of forward-facing cameras.
The trajectories estimated by our system for the 11 training sequences are shown in Figure 13 for the LO-only module and for the SLAM system (including loop closure), compared with ground truth. Quantitative results are shown in Table 3, where a large number of alternative methods are compared due to the popularity of this particular dataset. Note that two configurations of SiMpLE (Bhandari et al. (2024)) have been included (“online” and “offline”), with one of them being faster and the other more accurate. However, note that none would be able to run at sensor rate (10 Hz). Excepting IMLS-SLAM and SiMpLE, all others are fast enough to run in real-time at the sensor rate or faster.
@@
Figure 14: Estimated trajectories for the proposed LiDAR odometry system (”MOLA-LO”) applied to the KITTI-360 odometry dataset, compared to ground truth. Trajectories denoted as “MOLA-LO + LC” include the post-processing loop-closure stage. See discussion in §4.3.
Regarding the accuracy of the estimated trajectories for LO methods, it can be seen in Table 3 that the slower and more precise configuration of SiMpLE is slightly better than the rest, with KISS-ICP following next, then our LO system, then IMLS-SLAM. However, performance metrics for these SOTA methods are all quite similar, as demonstrated by an almost perfect tie in the average RTE metric used to evaluate all training ($\text{RTE}\sim 0.55\%$) and all testing sequences ($\text{RTE}\sim 0.61-0.62\%$).
Note that all shown results for KISS-ICP, SiMpLE, and our system, have been run using the same MOLA KITTI dataset source (see §3.2), which includes the $0.205^{\circ}$ vertical angle correction of all original dataset scans due to a miscalibration; see Deschaud (2018). Another well-known issue with KITTI is the wrong ground truth in part of sequence 08, which explains the relative elevated apparent errors of all methods in that sequence.
@@ 4.3 KITTI-360 dataset
This automotive dataset is a follow-up of the first KITTI dataset, including more sensors (fish-eye cameras, 2D LiDAR, and IMU) and longer driving sequences in suburban scenarios (Liao et al. (2022)).
@@
Figure 15: Estimated trajectories for the proposed LiDAR odometry system (”MOLA-LO”) applied to the Paris LuCo dataset, compared to KISS-ICP, and to ground truth. Trajectories denoted as “MOLA-LO + LC” include loop-closure. See discussion in §4.4.
Trajectories estimated by our system are shown in Figure 14, while quantitative performance results, compared to KISS-ICP as SOTA baseline, are summarized in Table 4. As can be seen in the table, where we evaluated the ATE metric, our system achieves better accuracy than KISS-ICP in 6 out of 8 sequences. Regarding our SLAM method with loop closures, and given that most KITTI-360 sequences, except for 03, 07, and 10, have multiple loop-closures, obtained ATE is further reduced. This dataset also includes 4 evaluation sequences without public ground truth for benchmarking on the “trajectory estimation” public leaderboard website [^6] See: https://www.cvlibs.net/datasets/kitti-360/leaderboard_semantic_slam.php?task=trajectory., where our SLAM solution ranks 3rd out of 5 submissions at the time of writing. In this leaderboard, more modern than the original KITTI, the metric that determines the classification is APE, and the results available online are: CT-ICP2 (LiDAR SLAM, Dellenbach et al. (2022)) (0.50 m), SOFT2 (stereo visual SLAM, Cvišić et al. (2022)) (0.70 m), our SLAM system (0.72 m), ORB-SLAM2 (stereo visual SLAM, Mur-Artal and Tard’os (2017)) (1.92 m), and SUMA++ (LiDAR SLAM, Chen et al. (2019)) (3.13 m).
@@ 4.5 NTU VIRAL: two 3D LiDARs
From the accuracy metrics, it is clear that methods using inertial sensors (LIOSAM and VIRAL SLAM) have the best performance, although we believe that the smarter sampling of LiDAR scans using edge and plane key points (Nguyen et al. (2021b)) plays a more significant role in achieving better accuracy than using an IMU; this topic would require further research. In contrast, LO methods KISS-ICP and ours, using only one (horizontal) LiDAR, have trajectory errors at least one order of magnitude larger than the aforementioned LIO methods, with KISS-ICP performing poorer in this case. Analyzing the spatial trajectories, it becomes clear that vertical movements of the drone are where these two methods perform worst, which can be explained by the simplistic uniform sampling of point clouds (instead of searching for specific features), which makes hard to distinguish vertical motions from remaining static. With this insight in mind, we provide an example of the flexibility of the present framework by adding two additional experiments evaluating MOLA-LO (the last two rows in Table 6). First, we put at test the capability of our system to handle several LiDARs, by enabling both vertical and horizontal LiDARs as inputs, with the idea of helping determining vertical motions. Our system (recall Figure 3 and the default pipeline in Figure 6) then waits for two consecutive scans before populating the observation metric map, taking into account the time offset between both scans to accurately de-skew both in a spatially coherent way. The rest remains exactly the same for the “MOLA-LO (2 LiDARs)” experiment shown in the table. As can be seen, using this vertical scanner indeed improves accuracy, but it is still far from LIO techniques. Note that all these results for MOLA-LO use the same configuration than in all other automotive datasets, with an automatic determination of the minimum points range to be filtered to prevent observations of the vehicle body. In this particular dataset, the smaller size of the drone makes that these rules actually prevent using valid information about nearby objects, but we did not manually tune this filter to ensure all datasets are run with the same configuration. Then, in a second experiment we modified the default pipeline in Figure 6 to create two local maps: “near map” and “far map”, for nearby and distant points, respectively. ICP pipeline blocks are then configured to search for potential pairings between nearby LiDAR scan points and the “near map”, and between all LiDAR points and the “far map” (interested readers can check the details in the configuration YAML file lidar3d-near-far.yaml available online). This configuration corresponds to the last row in the table. As can be seen, this method now has errors in the order of magnitude of LIO techniques, even surpassing LIOSAM on some particular sequences. Finally, using a smarter sampling of feature points would probably further improve the results, but this point is left for future works.
@@ 4.6 HILTI 2021 dataset
This dataset, presented in Helmberger et al. (2022), comprises several challenging sequences indoors and mixed indoor-outdoor scenarios recorded with an Ouster OS0-64 LiDAR. From the published sequences, only two of them have full SE(3) ground truth. From those two, we selected the largest sequence (the Drone Testing Area) for benchmarking our LO system. The resulting trajectories are illustrated in Figure 16. Quantitatively, we measured the rmse ATE for MOLA-LO against ground truth obtaining 0.18 m, slightly better than KISS-ICP (Vizzo et al. (2023)) which obtains an ATE of 0.21 m. It is noteworthy that our system fails to converge for some of the other sequences in this dataset (whose ground truth is not publicly released), where feature-extraction or multi-modality seem essential for robustly cope with scenarios like narrow, featureless indoor spaces.
@@
Figure 17: Estimated trajectory by our LO system for the VoxGraph dataset, together with the (partial) ground truth. See discussion in §4.7
@@ 4.7 Voxgraph dataset
This aerial dataset, presented in Reijgwart et al. (2019), comprises one sequence of a drone flight outdoors, equipped with an Ouster OS1-64 LiDAR and featuring RTK-based ground truth. We compare the ATE metric achieved by several LO SOTA methods in Table 7, where it can be seen that both KISS-ICP and ours achieve a significant better accuracy, with KISS-ICP having a small advantage in this case. The estimated trajectory for our method and the partially-available ground truth are illustrated in Figure 17. Note that ATE metrics have been only evaluated for those segments of the trajectory with corresponding ground truth. As in the HILTI Drone Testing Arena sequence (§4.6), loop closure is not required due to the small size of the environment, hence we only evaluate LO methods.
@@ 4.8 DARPA Subterranean dataset
Since loop closures are not relevant in these scenarios, only LO methods were evaluated. Two SOTA representative methods, KISS-ICP (Vizzo et al. (2023)) and SiMpLE (Bhandari et al. (2024)), have been run side by side with our system by using MOLA wrapper modules to ensure identical conditions for input data. KISS-ICP was run with its default parameters, just like system, run with its default configuration and pipelines (§3.8). In turn, for SiMpLE we used its slowest and most accurate configuration (“offline”), already evaluted for KITTI in §4.2.
@@
Figure 18: Estimated trajectories for the proposed LiDAR odometry system (”MOLA-LO”) applied to the DARPA Subterranean Final Event dataset, compared to ground truth. See discussion in §4.8.
@@
Figure 19: Estimated trajectories for the proposed LiDAR odometry system (”MOLA-LO”) applied to the NewerCollege dataset, compared to ground truth. Trajectories denoted as “MOLA-LO + LC” include the post-processing loop-closure stage discussed in section 3.13. Continues in Figure 20. See discussion in §4.9.
@@
Figure 20: Continuation of Figure 19 with more results for the Newer College dataset. See discussion in §4.9.
Table 9: Absolute translational error (ATE) (rmse values in meters as reported by “evo_ape -a”) for the NewerCollege dataset. Bold means best accuracy in each sequence and category. Divergence is shown as $\times$.
@@
0.68 m 0.54 m 0.12 m 0.06 m 0.12 m 0.11 m 3.85 m 0.08 m 0.08 m 0.12 m 0.07 m 0.08 m 0.09 m 0.11 m
@@
Figure 21: Detailed view of the Maths Hard (MH) sequence (Newer College dataset) in Figure 20. Observe the highly dynamic rotations imposed to the sensor in all six degrees of freedom with the intention of making this sequence a challenge for LO methods, and how our method keeps its estimation close to ground truth at all times. See discussion in §4.9.

### v2

@@ 0.81 $d/hm$
MOLA-LO
Config.: Always update map
2.11%
0.65 $d/hm$
10.58 m
2.09%
0.69 $d/hm$
9.65 m
2.33%
0.71 $d/hm$
10.18 m
2.53%
0.63 $d/hm$
10.88 m
2.27%
0.61 $d/hm$
9.18 m
1.87%
0.66 $d/hm$
9.14 m
3.25%
0.71 $d/hm$
35.13 m
3.08%
0.65 $d/hm$
31.25 m
2.50%
0.52 $d/hm$
29.55 m
4.06%
0.58 $d/hm$
466.81 m
4.58%
0.66 $d/hm$
512.25 m
4.85%
0.81 $d/hm$
413.81 m
24 ms
MOLA-LO
Config.: 3D-NDT
2.09%
0.64 $d/hm$
8.77 m
2.05%
0.68 $d/hm$
7.44 m
2.30%
0.71 $d/hm$
7.79 m
2.53%
0.61 $d/hm$
9.20 m
2.18%
0.61 $d/hm$
6.74 m
1.89%
0.65 $d/hm$
6.83 m
3.19%
0.69 $d/hm$
24.91 m
3.08%
0.66 $d/hm$
21.51 m
2.49%
0.50 $d/hm$
20.30 m
3.91%
0.56 $d/hm$
317.56 m
4.45%
0.63 $d/hm$
337.94 m
4.68%
0.78 $d/hm$
199.25 m
39 ms
MOLA-LO + LC (ours)
(default LO + GNSS)
@@ 0.74 $d/hm$
MOLA-LO + LC (ours)
(default LO, no GNSS)
2.08%
0.63 $d/hm$
2.57 m
1.98%
0.66 $d/hm$
2.56 m
2.34%
0.72 $d/hm$
2.76 m
2.50%
0.65 $d/hm$
5.31 m
2.05%
0.58 $d/hm$
2.75 m
1.76%
0.61 $d/hm$
1.91 m
3.79%
0.87 $d/hm$
22.76 m
3.21%
0.65 $d/hm$
9.09 m
2.52%
0.49 $d/hm$
6.93 m
4.08%
0.57 $d/hm$
381.54 m
4.61%
0.65 $d/hm$
419.82 m
4.80%
0.80 $d/hm$
315.43 m
42 ms
Figure 14: Summary of absolute translational error (ATE) for the MulRan dataset.
@@ 0.80 $d/hm$
Figure 15: Relative translational error (RTE) (%) for the training sequences in the KITTI visual-LiDAR odometry dataset.
@@ 4.1 MulRan dataset
This multimodal dataset comprises Ouster OS1-64 LiDAR data, consumer-grade GNSS, and precise ground truth for 12 driving sequences (Kim et al., 2020). All sequences include several loop closures, except the longest ones (Sejong{01,02,03}) which only include one, although the longest loop in all the benchmarked datasets ($>20$ km).
We have compared our LO system in two configurations, (i) its default one and (ii) its 3D-NDT version, against two other state-of-the-art LO methods: KISS-ICP (Vizzo et al., 2023) using its default settings, and SiMpLE (Bhandari et al., 2024) using their parameter set optimized for this dataset. Note that only SiMpLE has parameters optimized for this particular dataset, while our two configurations and KISS-ICP use the same parameters for all datasets. MOLA wrappers have been provided for both methods to provide a fair comparison in terms of identical API to provide input raw data and to collect estimated trajectories. The MOLA dataset source (see Section 3.13) for MulRan takes care of removing LiDAR scans at the beginning or end of each sequence that have no associated ground truth. In this way, we can use the KITTI evaluation metrics (RTE and RRE) apart of APE (see metrics definitions in Section 4). Consumer-grade GNSS observations are not used by our LO system, but they are while applying the loop-closure post-processing stage (explained in Section 3.12).
@@
Figure 16: Estimated trajectories for the proposed LiDAR odometry system (”MOLA-LO”) applied to the KITTI odometry dataset, compared to ground truth. Trajectories denoted as “MOLA-LO + LC” include loop-closure. See discussion in Section 4.2.
Sample trajectories from our method are shown in Figure 13, compared to GT. It can be seen that the output from loop closure is nearly indistinguishable from GT. A quantitative benchmark is provided in Table 2, where all three metrics (RTE, RRE, APE) are provided, together with average run times per LiDAR scan (measured on an Intel i7-8700 at 3.20GHz). A graphical summary of ATE metrics is also provided in Figure 14. All LO methods are faster than the sensor rate (10 Hz), hence all of them would be able to run onboard in real-time. An interesting observation is that there is not always a best performing method for a given sequence, since different methods may have the lowest values for one given metric only: LO methods would rank in a different order depending on the metric of interest. That said, for this dataset we can see how our 3D-NDT configuration clearly outperforms the others, despite not being specifically tuned for this dataset as SiMpLE is. Our SLAM version (reflected as “MOLA LO+LC” in the table) achieves, obviously, better results than the pure LO methods, hence their values are not marked in bold since the method is qualitatively different.
@@ 4.2 KITTI-odometry dataset
The KITTI odometry dataset (Geiger et al., 2013) was the first widely spread benchmark for LiDAR and visual odometry and SLAM methods, hence its popularity and importance. It comprises 11 training sequences (numbered 00-10) with public ground truth and other 11 evaluation sequences with undisclosed ground truth trajectories. Available sensors in this driving dataset include 3D LiDAR (a Velodyne HDL-64E) and two pairs of forward-facing cameras.
The trajectories estimated by our system (with its default configuration) for the 11 training sequences are shown in Figure 16 for the LO-only module and for the SLAM system (including loop closure), compared with ground truth. Quantitative results are shown in Table 3 and graphically in Figure 15. A larger number of alternative methods are compared in this case due to the popularity of this particular dataset. Note that two configurations of SiMpLE (Bhandari et al., 2024) have been included (“online” and “offline”), with one of them being faster and the other more accurate. However, note that none would be able to run at sensor rate (10 Hz). Excepting IMLS-SLAM and SiMpLE, all others are fast enough to run in real-time at the sensor rate or faster.
@@
Figure 17: Estimated trajectories for the proposed LiDAR odometry system (”MOLA-LO”) applied to the KITTI-360 odometry dataset, compared to ground truth. Trajectories denoted as “MOLA-LO + LC” include the post-processing loop-closure stage. See discussion in Section 4.3.
Regarding the accuracy of the estimated trajectories for LO methods, it can be seen in Table 3 that the slower and more precise configuration of SiMpLE is slightly better than the rest, although accuracy metrics for these SOTA methods are all quite similar, as demonstrated by an almost perfect tie in the average RTE metric used to evaluate all training ($\text{RTE}\sim 0.55\%$) and all testing sequences ($\text{RTE}\sim 0.61-0.62\%$).
Note that all shown results for KISS-ICP, SiMpLE, and our system, have been run using the same MOLA KITTI dataset source (see Section 3.13), which includes the $0.205^{\circ}$ vertical angle correction of all original dataset scans due to a miscalibration; see Deschaud (2018). Another well-known issue with KITTI is the wrong ground truth in part of sequence 08, which explains the relative elevated apparent errors of all methods in that sequence.
@@ 4.3 KITTI-360 dataset
This automotive dataset is a follow-up of the first KITTI dataset, including more sensors (fish-eye cameras, 2D LiDAR, and IMU) and longer driving sequences in suburban scenarios (Liao et al., 2022).
@@
Figure 18: Estimated trajectories for the proposed LiDAR odometry system (”MOLA-LO”) applied to the Paris LuCo dataset, compared to KISS-ICP, and to ground truth. Trajectories denoted as “MOLA-LO + LC” include loop-closure. See discussion in Section 4.4.
Trajectories estimated by our system are shown in Figure 17, while quantitative performance results, compared to KISS-ICP as SOTA baseline, are summarized in Table 4. As can be seen in the table, where we evaluated the ATE metric, our system achieves better accuracy than KISS-ICP in 6 out of 8 sequences. Regarding our SLAM method with loop closures, and given that most KITTI-360 sequences, except for 03, 07, and 10, have multiple loop-closures, obtained ATE is further reduced. This dataset also includes 4 evaluation sequences without public ground truth for benchmarking on the “trajectory estimation” public leaderboard website [^5] See: https://www.cvlibs.net/datasets/kitti-360/leaderboard_semantic_slam.php?task=trajectory., where our SLAM solution ranks 3rd out of 5 submissions at the time of writing. In this leaderboard, more modern than the original KITTI, the metric that determines the classification is APE, and the results available online are: CT-ICP2 with LiDAR SLAM in (Dellenbach et al., 2022) (0.50 m), SOFT2 with stereo visual SLAM in (Cvišić et al., 2022) (0.70 m), our SLAM system (0.72 m), ORB-SLAM2 with stereo visual SLAM in (Mur-Artal and Tardos, 2017) (1.92 m), and SUMA++ with LiDAR SLAM in (Chen et al., 2019) (3.13 m).
@@ 4.5 NTU VIRAL: two 3D LiDARs
From the accuracy metrics, it is clear that methods using inertial sensors (LIOSAM and VIRAL SLAM) have the best performance, although we believe that the smarter sampling of LiDAR scans using edge and plane key points (Nguyen et al., 2021b) plays a more significant role in achieving better accuracy than using an IMU; this topic would require further research. In contrast, LO methods KISS-ICP and ours, using only one (horizontal) LiDAR, have trajectory errors at least one order of magnitude larger than the aforementioned LIO methods, with KISS-ICP performing poorer in this case. Analyzing the spatial trajectories, it becomes clear that vertical movements of the drone are where these two methods perform worst, which can be explained by the simplistic uniform sampling of point clouds (instead of searching for specific features), which makes hard to distinguish vertical motions from remaining static. With this insight in mind, we provide an example of the flexibility of the present framework by adding two additional experiments evaluating MOLA-LO (the last two rows in Table 6). First, we put at test the capability of our system to handle several LiDARs, by enabling both vertical and horizontal LiDARs as inputs, with the idea of helping determining vertical motions. Our system (recall Figure 3 and the default pipeline in Figure 6) then waits for two consecutive scans before populating the observation metric map, taking into account the time offset between both scans to accurately de-skew both in a spatially coherent way. The rest remains exactly the same for the “MOLA-LO (2 LiDARs)” experiment shown in the table. As can be seen, using this vertical scanner indeed improves accuracy, but it is still far from LIO techniques. Note that all these results for MOLA-LO use the same configuration than in all other automotive datasets, with an automatic determination of the minimum points range to be filtered to prevent observations of the vehicle body. In this particular dataset, the smaller size of the drone makes that these rules actually prevent using valid information about nearby objects, but we did not manually tune this filter to ensure all datasets are run with the same configuration. Then, in a second experiment we modified the default pipeline in Figure 6 to create two local maps: “near map” and “far map”, for nearby and distant points, respectively. ICP pipeline blocks are then configured to search for potential pairings between nearby LiDAR scan points and the “near map”, and between all LiDAR points and the “far map” (interested readers can check the details in the configuration YAML file lidar3d-near-far.yaml available online). This configuration corresponds to the last row in the table. As can be seen, this method now has errors in the order of magnitude of LIO techniques, even surpassing LIOSAM on some particular sequences. Finally, using a smarter sampling of feature points would probably further improve the results, but this point is left for future works.
@@ 4.6 HILTI 2021 dataset
This dataset, presented in Helmberger et al. (2022), comprises several challenging sequences indoors and mixed indoor-outdoor scenarios recorded with an Ouster OS0-64 LiDAR. From the published sequences, only two of them have full SE(3) ground truth. From those two, we selected the largest sequence (the Drone Testing Area) for benchmarking our LO system. The resulting trajectories are illustrated in Figure 19. Quantitatively, we measured the rmse ATE for MOLA-LO against ground truth obtaining 0.18 m, slightly better than KISS-ICP (Vizzo et al., 2023) which obtains an ATE of 0.21 m. It is noteworthy that our system fails to converge for some of the other sequences in this dataset (whose ground truth is not publicly released), where feature-extraction or multi-modality seem essential for robustly cope with scenarios like narrow, featureless indoor spaces.
@@
MOLA-LO (3D-NDT)
0.250 m
Figure 20: Estimated trajectory by our LO system for the VoxGraph dataset, together with the (partial) ground truth. See discussion in Section 4.7
@@ 4.7 Voxgraph dataset
This aerial dataset, presented in Reijgwart et al. (2019), comprises one sequence of a drone flight outdoors, equipped with an Ouster OS1-64 LiDAR and featuring RTK-based ground truth. We compare the ATE metric achieved by several LO SOTA methods in Table 7, where it can be seen that both KISS-ICP and ours achieve a significant better accuracy, with KISS-ICP having a small advantage in this case. The estimated trajectory for our method and the partially-available ground truth are illustrated in Figure 20. Note that ATE metrics have been only evaluated for those segments of the trajectory with corresponding ground truth. As in the HILTI Drone Testing Arena sequence (Section 4.6), loop closure is not required due to the small size of the environment, hence we only evaluate LO methods.
@@ 4.8 DARPA Subterranean dataset
Since loop closures are not relevant in these scenarios, only LO methods were evaluated. Two SOTA representative methods, KISS-ICP (Vizzo et al., 2023) and SiMpLE (Bhandari et al., 2024), have been run side by side with our system by using MOLA wrapper modules to ensure identical conditions for input data. KISS-ICP was run with its default parameters, just like system, run with its default configuration and pipelines (Section 3.7). In turn, for SiMpLE we used its slowest and most accurate configuration (“offline”), already evaluted for KITTI in Section 4.2.
@@
Figure 21: Estimated trajectories for the proposed LiDAR odometry system (”MOLA-LO”) applied to the DARPA Subterranean Final Event dataset, compared to ground truth. See discussion in Section 4.8.
@@
Figure 22: Estimated trajectories for the proposed LiDAR odometry system (”MOLA-LO”) applied to the NewerCollege dataset, compared to ground truth. Trajectories denoted as “MOLA-LO + LC” include the post-processing loop-closure stage discussed in section 3.12. Continues in Figure 23. See discussion in Section 4.9.
@@
Figure 23: Continuation of Figure 22 with more results for the Newer College dataset. See discussion in Section 4.9.
Table 9: Absolute translational error (ATE) (rmse values in meters as reported by “evo_ape -a”) for the NewerCollege dataset. Bold means best accuracy in each sequence and category. Divergence is shown as $\times$. The version of MOLA-LO that always updates the local map is discussed in Section 7.2.
@@
0.68 m 0.54 m 0.12 m 0.06 m 0.12 m 0.11 m 0.90 m 0.08 m 0.08 m 0.12 m 0.07 m 0.08 m 0.09 m 0.11 m
MOLA-LO
(Always updates
local map)
$\times$
(5.11 m)
$\times$
(84.37 m)
0.39 m 0.09 m
$\times$
(4.93 m)
0.78 m 1.38 m 0.08 m 2.99 m 0.10 m 0.08 m 0.46 m
$\times$
(8.06 m)
0.34 m
@@
Figure 24: Detailed view of the Maths Hard (MH) sequence (Newer College dataset) in Figure 23. Observe the highly dynamic rotations imposed to the sensor in all six degrees of freedom with the intention of making this sequence a challenge for LO methods, and how our method keeps its estimation close to ground truth at all times. See discussion in Section 4.9.